OpenStudy (anonymous):

Medal for correct answer!!!! Write the equation of the horizontal line passing through the point (-1, -8).

OpenStudy (anonymous):

horizontal line looks like \(y=number\) in your case the number is \(-8\)

OpenStudy (anonymous):

whats the EQUATION??

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

therefore the line cannot be horizontal

OpenStudy (anonymous):

DO IT FOR ME FOR A MEDALLLL!!!

OpenStudy (anonymous):

A horizontal line that PASSES THROUGH...

OpenStudy (anonymous):

it can't be horizontal if it passes through those points as they are formatted here as (x, y)

OpenStudy (anonymous):

horizontal means that it goes from side to side

OpenStudy (anonymous):

y=0+(-8)

OpenStudy (anonymous):

y=b

OpenStudy (anonymous):

there's your equation hahaha B^]

OpenStudy (anonymous):

y=((0(-1))+(-8))

OpenStudy (anonymous):

(0*-1)=0 +(-8)=(-8)

OpenStudy (anonymous):

\[y =mx+b\]

OpenStudy (anonymous):

\[-8=(-1m+(-8))\]

OpenStudy (anonymous):

ahem.... I believe a medal is in orderrrr